Two cars leave at the same time traveling in different directions. One goes North at 6 mph and the other goes west at 8mph. How far apart are they after 4 hrs.
(6*4)^2+(8*4)^2=d^2
24^2+32*2=d^2
576+1024=d^2
1600=d^2
d=sqrt1600
d=40 miles is the distance between two cars after 4 hours

Where did the (2) come from in this answer?

Answer by KnightOwlTutor(293)

You are using the Pythagorean theorem a^2 +b^2=c^2
The the directions that the cars are traveling are creating a right angle
North is up and West is left. The resultant hypotenuse is distance between the two cars.
We know that after 4 hours car that travels 6mph has gone a distance of 4*6=24miles in the North direction
After 4 hours he car that travels 8mph has traveled a distane of 8*4=32 miles in the West direction
Each distance corresponds to a side of the triangle that meet in a right angle
Therefore 24^2+32^2=d^2
576+1024 =d^2
1600=d^2
take the square root of each side
d=40 miles

d=40 miles is the distance between two cars after 4 hours